# Robust Optimization with Diffusion Models for Green Security

### v1

@@ 1 Introduction
However, existing adversary predictive models either lack uncertainty quantification [Kar et al., 2017, Gurumurthy et al., 2018] or provide only parameterized predictive distributions with limited expressiveness [Xu et al., 2020b]. In reality, adversarial behaviors—such as those of poachers—are high-dimensional and highly complex, driven by diverse motivations, constraints, and strategies. Capturing the full extent of uncertainty is particularly challenging in the strategic environments, where conventional models may struggle to account for the variability of real-world threats.
@@ 2 Related Works
Diffusion Models Diffusion models have achieved remarkable success across various generative modeling tasks, including image generation [Song et al., 2021, Ho et al., 2020], optimization [Kong et al., 2024], and scientific discovery [Gruver et al., 2024, Watson et al., 2023, Kong et al., 2023]. These models are particularly adept at capturing complex, high-dimensional distributions, making them a powerful tool for diverse applications. Conditional diffusion models extend this capability by integrating contextual information to guide the generative process. By conditioning on textual descriptions, semantic masks, or other relevant features, these models enable tasks such as text-to-image generation [Saharia et al., 2022], image-to-image translation [Saharia et al., 2021], and time series forecasting [Shen and Kwok, 2023].
Double Oracle for Robust Optimization Prior work has framed robust optimization as a two-player zero-sum game Mastin et al. [2015], Gilbert and Spanjaard [2017], where the optimizing player selects a potentially randomized feasible strategy, while an adversary chooses problem parameters to maximize regret. The double oracle (DO) algorithm is a standard method for computing equilibria in such games  McMahan et al. [2003], Adam et al. [2021] and has been applied to robust influence maximization in social networks Wilder et al. [2017], robust patrol planning Xu et al. [2021], robust submodular optimization Wilder [2018], and robust policy design for restless bandits Killian et al. [2022]. However, these applications restrict the uncertainty set to a compact interval. In contrast, our problem involves a diffusion model that provides full distribution-level predictions, making the uncertainty set a space of distributions, which introduces new theoretical challenges in applying double oracle.
@@ 3 Preliminaries on Diffusion Model
Conditional Extension. This diffusion framework can be naturally extended to include additional context $\mathbf{c}$. In a conditional diffusion model, the score network becomes $s_{\theta}(\mathbf{z}^{t},t,\mathbf{c})$, so that at each step the denoising is informed by side information such as class labels, textual descriptions, or other relevant features. This conditional approach enables the generation of samples that match not only the learned data distribution but also the specific context $\mathbf{c}$, making it particularly useful for tasks in which external conditions strongly influence the underlying data generation process.
@@ 4 Problem Formulation
Table 1: Forecasting performance on poaching data.
@@ 5.3 Sampling with Twisted Sequential Monte Carlo
Under regularity conditions on the score function, as the number of particles $N\to\infty$, we have
@@ 5.4 Convergence Analysis
Proof. See Appendix G.
@@ 7 Conclusion
Appendix for DiffOracle
@@
Appendix A Algorithm 2
The full procedure of the twised Sequential Monte-Carlo sampler is given in Algorithm 2.
Algorithm 2 Twisted SMC for Diffusion Model
3:for $t=T,\dots,1$ do
4:    Resample:
6:    for $k=1\dots K$ do
10:    end for
11:end for
12:Output: Weighted particles $\{\mathbf{z}_{k}^{0},\,w_{k}^{0}\}_{k=1}^{K}$

### v2

@@ 1 Introduction
However, existing adversary predictive models [Kar et al., 2017, Gurumurthy et al., 2018] either lack uncertainty quantification [Kong et al., 2023b, Li et al., 2023, Kong et al., 2024] or provide only parameterized predictive distributions with limited expressiveness [Xu et al., 2020b]. In reality, adversarial behaviors—such as those of poachers—are high-dimensional and highly complex, driven by diverse motivations, constraints, and strategies. Capturing the full extent of uncertainty is particularly challenging in the strategic environments, where conventional models may struggle to account for the variability of real-world threats.
@@ 2 Related Works
Diffusion Models Diffusion models have achieved remarkable success across various generative modeling tasks, including image generation [Song et al., 2021, Ho et al., 2020], decision-making [Kong et al., 2025a, b], and scientific discovery [Gruver et al., 2024, Watson et al., 2023, Kong et al., 2023a]. These models are particularly adept at capturing complex, high-dimensional distributions, making them a powerful tool for diverse applications. Conditional diffusion models extend this capability by integrating contextual information to guide the generative process. By conditioning on textual descriptions, semantic masks, or other relevant features, these models enable tasks such as text-to-image generation [Saharia et al., 2022], image-to-image translation [Saharia et al., 2021], and time series forecasting [Shen and Kwok, 2023].
Double Oracle for Robust Optimization Prior work has framed robust optimization as a two-player zero-sum game [Mastin et al., 2015, Gilbert and Spanjaard, 2017], where the optimizing player selects a potentially randomized feasible strategy, while an adversary chooses problem parameters to maximize regret. The double oracle (DO) algorithm is a standard method for computing equilibria in such games  [McMahan et al., 2003, Adam et al., 2021] and has been applied to robust influence maximization in social networks [Wilder et al., 2017], robust patrol planning [Xu et al., 2021], robust submodular optimization [Wilder, 2018], and robust policy design for restless bandits [Killian et al., 2022]. However, these applications restrict the uncertainty set to a compact interval. In contrast, our problem involves a diffusion model that provides full distribution-level predictions, making the uncertainty set a space of distributions, which introduces new theoretical challenges in applying double oracle.
@@ 3 Preliminaries on Diffusion Model
Conditional Extension. This diffusion framework can be naturally extended to include additional context $\mathbf{c}$. In a conditional diffusion model [Ho and Salimans, 2021], the score network becomes $s_{\theta}(\mathbf{z}^{t},t,\mathbf{c})$, so that at each step the denoising is informed by side information such as class labels, textual descriptions, or other relevant features. This conditional approach enables the generation of samples that match not only the learned data distribution but also the specific context $\mathbf{c}$, making it particularly useful for tasks in which external conditions strongly influence the underlying data generation process.
@@ 4 Problem Formulation
Table 1: Forecasting performance in terms of mean squared error (MSE) on poaching data.
@@ 5.3 Sampling with Twisted Sequential Monte Carlo
Algorithm 2 Twisted SMC for Diffusion Model
3:for $t=T,\dots,1$ do
4:    Resample:
6:    for $k=1\dots K$ do
10:    end for
11:end for
12:Output: Weighted particles $\{\mathbf{z}_{k}^{0},\,w_{k}^{0}\}_{k=1}^{K}$
@@
(Informal) Under regularity conditions on the score function, as the number of particles $N\to\infty$, we have
@@ 5.4 Convergence Analysis
Proof.
We provide a sketch of the proof here and defer the full details to Appendix G. The key steps for proving Item 1 are as follows:
  • Step 1: We bound the utility estimation error for any mixed strategy pair at iteration $i$ by the maximum estimation error over all entries in the payoff matrix.
  • Step 2: We show that, under our finite sampling scheme, the probability that the maximum cell-wise error exceeds $\epsilon/4$ is nonzero only during the first $i_{r}$ iterations, for some finite $i_{r}$.
  • Step 3: We treat the strategies generated in the first $i_{r}$ rounds as the initial strategy set in the standard Double Oracle (DO) algorithm [Adam et al., 2021]. We then adapt the original convergence proof to account for the error introduced by finite sampling, which is now bounded by $\epsilon/4$.
By relaxing the error bound in Item 1, we obtain convergence within a finite number of iterations. The additional approximation error in Item 2 stems from two sources: (1) enforcing finite termination, and (2) using estimated utilities of mixed strategy pairs when evaluating the stopping condition.
@@ 7 Conclusion
Acknowledgements
We are thankful to the Uganda Wildlife Authority for granting us access to incident data from Murchison Falls National Park. We also thank Charles Emogor for the insightful discussions and the anonymous reviewers for their valuable feedback. This work was supported by ONR MURI N00014-24-1-2742.
Appendix for Robust Optimization with Diffusion Models for Green Security
@@ Acknowledgements
Appendix A More Details on Diffusion Models
[…]
Conditional Extension. This diffusion framework can be naturally extended to include additional context $\mathbf{c}$. In a conditional diffusion model [Ho and Salimans, 2021], the score network becomes $s_{\theta}(\mathbf{z}^{t},t,\mathbf{c})$, so that at each step the denoising is informed by side information such as class labels, textual descriptions, or other relevant features. This conditional approach enables the generation of samples that match not only the learned data distribution but also the specific context $\mathbf{c}$, making it particularly useful for tasks in which external conditions strongly influence the underlying data generation process.
Rather than directly estimating the score function $s_{\theta}(\mathbf{z}^{t},t)$, Denoising Diffusion Probabilistic Models (DDPM) [Ho et al., 2020] reformulate the learning objective as a noise prediction task. This reparameterization leverages the closed-form expression of the forward process:
$\mathbf{z}^{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$
where $\bar{\alpha}_{t}$ denotes the cumulative product of noise schedules. The training objective becomes recovering the noise $\boldsymbol{\epsilon}$ that perturbed $\mathbf{z}^{0}$ to form $\mathbf{z}^{t}$. A neural network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}^{t},t)$ is trained to approximate this noise, which corresponds to learning the score function up to a time-dependent scaling:
$s_{\theta}(\mathbf{z}^{t},t)=-\frac{\boldsymbol{\epsilon}_{\theta}(\mathbf{z}^{t},t)}{\sqrt{1-\bar{\alpha}_{t}}}.$
Training then reduces to minimizing a simple mean squared error (MSE) loss between the true and predicted noise:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{\mathbf{z}_{0},\,\boldsymbol{\epsilon},\,t}\!\bigl{[}\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\|^{2}\bigr{]}.$
By training this conditional diffusion model on historical poaching data—augmented with contextual features $\mathbf{c}$—we learn $p_{\theta}(\mathbf{z}\mid\mathbf{c})$, a powerful and expressive model of poacher behavior. This enables us to capture complex, multimodal patterns of attacker responses, thereby supporting the development of robust patrol strategies discussed earlier. One challenge of this task is that poaching data is often very noisy [Zhuang et al., 2023].